OpenStudy (anonymous):

Traveling at 55 miles per hour, how many minutes, rounded to the nearest whole number, does it take to drive 310 miles from Boston to Syracuse?

OpenStudy (cwrw238):

speed = distance / time 55 = 310 / x where x is the time in hours solve for x then convert hours to minutes

OpenStudy (cwrw238):

cross multiply 55x = 310 now divide both sides by 55 x = 310/55 hours when you;ve worked that out multiply by 60 to give the time in minutes.

OpenStudy (anonymous):

THANK YOU SOOOO MUCH

OpenStudy (ali2x2):

remember to close this question. :)

OpenStudy (jacob902):

1 hr = 60 minutes now we will set up a proportion 55/60 = 310/x --- 55 miles to 60 minutes = 310 miles to x minutes cross multiply (55)(x) = (60)(310) 55x = 18600 x = 18600/55 x = 338 2/11 to the nearest whole number is 338